# MedUnifier: Unifying Vision-and-Language Pre-training on Medical Data with Vision Generation Task using Discrete Visual Representations

### v2

@@ 3.4 Text-grounded image generation (TIG)
The VQ-VAE Van Den Oord et al. (2017) offers a significant advantage over other generative models due to its ability to explicitly learn discrete visual representations. This feature aligns closely with the image-text encoder, which similarly learns discrete representations from a dictionary of learnable embeddings. Motivated by this, we chose to adopt VQ-VAE as the image generator in our framework, which forms a different unified Med-VLP framework from other studies Zhou et al. (2022b); Chen et al. (2023a, b). Additionally, the image-text encoder generates two distinct types of visual features: one for abstract visual representations, denoted as $\boldsymbol{f}^{q}$, and another for fine-grained, local visual embeddings, $\boldsymbol{f}^{v}_{local}$. Consequently, the image-text encoder can be viewed as a powerful multi-modal encoder, akin to the image encoder in conventional VAEs Van Den Oord et al. (2017); Higgins et al. (2017). Inspired by the work of Razavi et al. (2019), we developed hierarchical vector quantizers and corresponding image decoders. Figure 2 further illustrates the proposed text-grounded image generator. Our TIG module is specifically designed to capture fine-grained details through recovering pixel-level information from hierarchical multi-modal representations, enabling the model to identify subtle visual details.
@@
The feature vector $\boldsymbol{f}^{q}\in\mathbb{R}^{L_{q}\times d_{q}}$ contains textual implications derived from the co-training image-text encoder, which we regard as the top latent representation, denoted $\boldsymbol{z}^{top}\in\mathbb{R}^{L_{q}\times d_{q}}$. In contrast, $\boldsymbol{f}^{v}_{local}\in\mathbb{R}^{L_{v}\times d_{v}}$ does not include textual information. Therefore, we concatenate $\boldsymbol{f}^{v}_{local}$ with the aggregated textual representation, $\boldsymbol{f}^{t}_{[CLS]}\in\mathbb{R}^{1\times d_{t}}$, along the feature dimension, resulting in the multi-modal bottom latent representation $\boldsymbol{z}^{bottom}\in\mathbb{R}^{L_{v}\times(d_{v}+d_{t})}$.
@@
where $\mathcal{Z}_{top}$ consisted of a nonlinear transformation, spatial positional encoding summer Wang and Liu (2019) and a residual block He et al. (2016). The illustration can be found in appendix 1. We reshaped $\boldsymbol{z}^{top}$ and then pass it through $\mathcal{Z}_{top}$, reaping $\boldsymbol{z}_{e}^{top}\in\mathbb{R}^{d_{top}\times h_{top}\times w_{top}}$. Followed by a vector quantization layer with a latent embedding space $e^{top}$, we gain discrete feature map $\boldsymbol{z}_{q}^{top}\in\mathbb{R}^{d_{top}\times h_{top}\times w_{top}}$:
@@
$\boldsymbol{z}_{e}^{bottom}$ $=\mathcal{Z}_{bottom}(\boldsymbol{z}^{bottom})$ (20)
$\boldsymbol{z}_{q}^{bottom}$ $=quantizer_{bottom}(\boldsymbol{z}_{e}^{bottom},\boldsymbol{z}_{q}^{top})$ (21)
where $\boldsymbol{z}^{bottom}$ are viewed as having spatial size $h_{bottom}\times w_{bottom}=\frac{H}{h}\times\frac{W}{w}$. Note that producing $\boldsymbol{z}_{q}^{bottom}$ is conditioning on $\boldsymbol{z}_{q}^{top}$.
We built hierarchical decoder $\mathcal{D}$ with deconvolutional layers to recover raw visual input from discrete multi-modal representations:
$\hat{x}^{i}=\mathcal{D}(\boldsymbol{z}_{q}^{top},\boldsymbol{z}_{q}^{bottom})$ (22)
@@
$\mathcal{L}_{tig}$ $=\frac{1}{N}\sum_{k=1}^{N}-\log p(x^{i}_{k}|\boldsymbol{z}_{q}^{top},\boldsymbol{z}_{q}^{bottom})$ (23)
@@ 3.5 Total learning objectives
We provide a comprehensive summary of all the VLP learning objectives and present the corresponding ultimate objective function:
@@
Four loss weights $\lambda$ were set to 1 in experiments. These weights were determined by ablation study (see appendix 5).
@@
ConVIRTZhang et al. (2022) 84.2 86.9 88.7 84.2 85.7 91.5 72.5 82.5 92.0
GLoRIAHuang et al. (2021) 84.1 86.8 89.1 85.1 88.5 92.1 66.5 80.5 88.0
BioViLBoecking et al. (2022) 82.0 85.4 88.6 79.8 81.6 90.5 - - -
LoVT Müller et al. (2022) 85.1 86.5 89.3 85.5 88.5 92.2 - - -
MGCA Wang et al. (2022a) 85.8 87.7 89.2 86.1 89.6 92.0 74.8 84.8 92.3
PRIOR Cheng et al. (2023) 85.7 87.1 89.2 87.2 89.1 92.3 - - -
@@
ConVIRTZhang et al. (2022) 46.5 53.9 53.8 20.0 45.4 35.5
GLoRIAHuang et al. (2021) 46.7 56.4 55.0 51.8 59.5 58.9
BioViLBoecking et al. (2022) 47.3 57.7 55.6 54.6 64.3 62.8
MedCLIP Wang et al. (2022d) 47.6 58.0 55.9 56.3 69.9 66.7
MGCA Wang et al. (2022a) 47.1 57.4 55.4 53.1 61.9 61.1
REFERS Zhou et al. (2022a) 52.4 59.9 58.6 60.6 71.9 69.0
CXR-CLIP You et al. (2023) 51.8 61.2 58.5 60.2 69.2 64.6
@@ Methods
ConVIRTZhang et al. (2022) 43.8 35.2 77.4
GLoRIAHuang et al. (2021) 47.5 45.0 68.3
BioViLBoecking et al. (2022) 48.5 42.2 77.1
MedCLIPWang et al. (2022d) 47.1 41.1 81.8
MGCA Wang et al. (2022a) 48.0 40.9 76.2
REFERS Zhou et al. (2022a) 49.5 41.8 78.0
CXR-CLIP You et al. (2023) 49.7 35.9 76.9
@@ 4 Experiments
We perform the pre-training on the current largest multi-modal medical dataset, MIMIC-CXR v2.0.0 Johnson et al. (2019) and evaluate our model on various downstream tasks, followed by an ablation study for probing purpose, which shows the proposed method’s superiority.
@@ 4.1 Implementation details
We employed a BERT model as the primary network for the image-text encoder and utilized ViT-g as the pre-trained ViT. The input image resolution was set to $224\times 224$, with a maximum text length of 95 tokens, and 32 learnable embeddings. The top and bottom codebook size were both set to 512 with feature dimension of 768. For optimization, we applied the AdamW optimizer Loshchilov (2017) with parameters $\beta_{1}=0.9$, $\beta_{2}=0.95$, and a weight decay of 0.05. A cosine learning rate decay schedule was used, with a peak learning rate of 1e-4. We incorporated a warm-up phase for the initial 5% of training steps, starting with a learning rate of 1e-5. The pre-training process was conducted on four NVIDIA A100 GPUs.The implementation details are further elaborated in appendix 2.
@@ 4.2 Medical Vision-and-Language Benchmark
To assess the effectiveness of our proposed method, we conducted experiments across three types of tasks: uni-modal, cross-modal, and multi-modal. We conducted experiments on MedUnifier with/without TIG loss and compared with previous studies as our main experimental results. All benchmark datasets used in experiments pertain specifically to radiology with detailed descriptions listed in the appendix 3. For state-of-the-art methods, we took reference from their papers and implemented them using their official codes: ConVIRTZhang et al. (2022), GLoRIAHuang et al. (2021), BioViLBoecking et al. (2022), LoVT Müller et al. (2022), MedCLIPWang et al. (2022d), MGCA Wang et al. (2022a), PRIOR Cheng et al. (2023), REFERS Zhou et al. (2022a), CXR-CLIP You et al. (2023), ViLTBannur et al. (2023), R2Gen Chen et al. (2020), PTUnifier Chen et al. (2023b), DCL Li et al. (2023c), MOTOR Lin et al. (2023), UniXGerLee et al. (2023a), RoentGenChambon et al. (2022), LLM-CXRLee et al. (2023b).
Uni-modal tasks     assess the learned visual representations for image modality by applying uni-modal masking within classification scope on datasets such as RSNA Pneumonia Shih et al. (2019), SIIM-ACR Zawacki et al. (2019), and COVIDx Wang et al. (2020). To examine the model’s data efficiency, we fine-tune it using different proportions of the training data (1%, 10%, or 100%).
@@
require models to effectively align both vision and language modalities. We conduct experiments across three subtasks: image-to-text retrieval (ITR), text-to-image retrieval (TIR), and zero-shot image classification (ZS). For the ITR and TIR tasks, we report cross-modal information retrieval metrics, specifically the mean Average Precision at K (mAP@K), using the MIMIC 5x200 dataset. In the ZS task, we employ MIMIC 5x200, CheXpert 5x200 Huang et al. (2021); Cheng et al. (2023), and draw 500 positive and 500 negative samples from the full RSNA Pneumonia dataset Shih et al. (2019) for evaluation purpose.
@@
generate uni-modal content through multi-modal interaction. We carry out two kinds of experiments including image-grounded text generation (ITG) and text-grounded image generation (TIG). For ITG, we adopt the MIMIC-CXR held-out test set to evaluate the quality of generated reports. Standard natural language generation (NLG) criteria are used to assess the performance, including BLEUn Papineni et al. (2002), METEOR Denkowski and Lavie (2011), and ROUGE-L Lin (2004). For TIG, following Razavi et al. (2019), we first train two Pixelsnail models Chen et al. (2018) to model multi-modal priors $\boldsymbol{z}_{q}^{top},\boldsymbol{z}_{q}^{bottom}$. Then we sample latent encodings from both priors and generate new medical images by decoding latent encodings. For quantitative analysis, we present FID Heusel et al. (2017) scores.
@@ 4.3 Results and Analyses
To validate the effectiveness of MedUnifier, we conduct experiments on the above vision-and-language benchmark. The results of the main experiments are presented in Table 1,2, 3, 4 5 and Figure 3. We observe several noteworthy findings in our results. First, our model outperforms prior studies on uni-modal tasks across various downstream datasets, as shown in Table 1. This improvement suggests that integrating TIG significantly enhances the model’s ability to learn more transferable visual representations. Second, for cross-modal retrieval, MedUnifier model achieves the highest performance, demonstrating a superior ability to understand and integrate cross-modal content compared to other models (see Table 2). The obtained results further highlight the better performance of MedUnifier, suggesting that our approach supports the necessary complementary semantic data for cross-modal retrieval. We also observe that excluding the TIG module slightly decreases MedUnifier’s effectiveness in text-image retrieval tasks, which may be a result of an over-reliance on ITG that leads to an imbalance in vision-language fusion. In addition, MedUnifier demonstrates better performance on zero-shot classification tasks for both the MIMIC 5x200 and RSNA datasets in Table 3. However, GLoRIA outperforms our model on the CheXpert 5x200 dataset, likely due to its pre-training on the full CheXpert dataset with accompanying medical reports. Nonetheless, the results indicate the effectiveness of employing prompt ensembles within the proposed method, leading to enhanced overall performance improvements. Table 4 demonstrates that our models, both with and without the TIG module, surpass previous methods for image-grounded medical report generation. The Med-VLP framework also gains substantial advantages from incorporating causal language modelling, which together contributes to its improved performance. Finally, we conduct both quantitative and qualitative analyses on text-grounded image generation tasks, as presented in Table 5 and Figure 3. It highlights that MedUnifier with TIG could achieve comparable performance for medical vision generation tasks with less model’s complexity (one-stage pre-training) then existing models Lee et al. (2023b, a); Chambon et al. (2022), all of which requires additional pre-training for image tokenizer or iterative denoising. From a direct visual inspection, the reconstructed visual samples are nearly indistinguishable from authentic radiographs. Moreover, synthetic samples generated from multi-modal priors demonstrate high diversity, highlighting their potential to augment out-of-distribution medical data effectively.
@@
DCLLi et al. (2023c) - - - 10.9 15.0 28.4
MOTORLin et al. (2023) - 21.3 - 15.6 19.3 31.4
ViLTBannur et al. (2023) - 21.3 - 9.2 - 29.6
R2Gen Chen et al. (2020) 35.3 21.8 14.5 10.3 14.2 27.7
PTUnifier Chen et al. (2023b) - - - 10.7 - -
@@
Methods FID $\downarrow$
UniXGenLee et al. (2023a) 78.2
RoentGenChambon et al. (2022) 42.4
LLM-CXRLee et al. (2023b) 22.8
Validation set 17.2
Reconstruction 27.2
MedUnifier 42.6
Table 5: FID score Heusel et al. (2017) on reconstructed and synthetic images, using the MIMIC-CXR pre-training dataset as the reference dataset. The FID score using MIMIC-CXR validation is also provided for a fair comparison. Reconstructed images are generated by passing the training set through our MedUnifier. Synthetic images are produced from sampled and decoded latent encodings using the trained Pixelsnail models and VAE decoder.
ID
Learning objectives
@@
ITR
ITC ITM ITG TIG mAP@1 mAP@5 mAP@10
1 53.7 63.8 59.9
2 55.2 64.1 60.3
3 57.4 65.4 60.3
4 58.5 65.2 60.8
5 60.7 66.6 61.7
Table 6: Ablation studies on the different types of objectives. The best performance is achieved using all learning objectives.

### v3

@@ 3.4 Text-grounded image generation (TIG)
The VQ-VAE [56] offers a significant advantage over other generative models due to its ability to explicitly learn discrete visual representations. This feature aligns closely with the image-text encoder, which similarly learns discrete representations from a dictionary of learnable embeddings. Motivated by this, we chose to adopt VQ-VAE as the image generator in our framework, which forms a different unified Med-VLP framework from other studies [72, 7, 10]. Additionally, the image-text encoder generates two distinct types of visual features: one for abstract visual representations, denoted as $\boldsymbol{f}^{q}$, and another for fine-grained, local visual embeddings, $\boldsymbol{f}^{v}_{local}$. Consequently, the image-text encoder can be viewed as a powerful multi-modal encoder, akin to the image encoder in conventional VAEs [56, 24]. Inspired by the work of [50], we developed hierarchical vector quantizers and image decoders. Figure 2 further illustrates the proposed text-grounded image generator. Our TIG module is specifically designed to capture fine-grained details through recovering pixel-level information from hierarchical multi-modal representations, enabling the model to identify subtle visual details.
@@
The features $\boldsymbol{f}^{q}\in\mathbb{R}^{L_{q}\times d_{q}}$ contains textual implications derived from the co-training image-text encoder, which we regard as the top latent representation, denoted $\boldsymbol{z}^{top}\in\mathbb{R}^{L_{q}\times d_{q}}$. In contrast, $\boldsymbol{f}^{v}_{local}\in\mathbb{R}^{L_{v}\times d_{v}}$ does not include textual information. Therefore, we concatenate $\boldsymbol{f}^{v}_{local}$ with the aggregated textual representation, $\boldsymbol{f}^{t}_{[CLS]}\in\mathbb{R}^{1\times d_{t}}$, along the feature dimension, resulting in the multi-modal bottom latent representation $\boldsymbol{z}^{bot}\in\mathbb{R}^{L_{v}\times(d_{v}+d_{t})}$.
@@
where $\mathcal{Z}_{top}$ consisted of a nonlinear transformation, spatial positional encoding summer [61] and a residual block [22]. The illustration can be found in appendix 7. We reshaped $\boldsymbol{z}^{top}$ and then pass it through $\mathcal{Z}_{top}$, reaping $\boldsymbol{z}_{e}^{top}\in\mathbb{R}^{d_{top}\times h_{top}\times w_{top}}$. Followed by a vector quantization layer with a latent embedding space $e^{top}$, we gain discrete feature map $\boldsymbol{z}_{q}^{top}\in\mathbb{R}^{d_{top}\times h_{top}\times w_{top}}$:
@@
$\boldsymbol{z}_{e}^{bot}$ $=\mathcal{Z}_{bot}(\boldsymbol{z}^{bot})$ (20)
$\boldsymbol{z}_{q}^{bot}$ $=quantizer_{bot}(\boldsymbol{z}_{e}^{bot},\boldsymbol{z}_{q}^{top})$ (21)
where $\boldsymbol{z}^{bot}$ are viewed as having spatial size $h_{bot}\times w_{bot}=\frac{H}{h}\times\frac{W}{w}$. Producing $\boldsymbol{z}_{q}^{bottom}$ is conditioning on $\boldsymbol{z}_{q}^{top}$.
We built hierarchical decoders $\mathcal{D}$ to recover raw images from discrete multi-modal representations:
$\hat{x}^{i}=\mathcal{D}(\boldsymbol{z}_{q}^{top},\boldsymbol{z}_{q}^{bot})$ (22)
@@
$\mathcal{L}_{tig}$ $=\frac{1}{N}\sum_{k=1}^{N}-\log p(x^{i}_{k}|\boldsymbol{z}_{q}^{top},\boldsymbol{z}_{q}^{bot})$ (23)
@@ 3.5 Total learning objectives
We provide a comprehensive summary of all the learning objectives and present the ultimate loss function:
@@
Four weights $\lambda$ were set to 1 in experiments. These weights were determined by ablation study (see appendix 11).
@@
ConVIRT[69] 84.2 86.9 88.7 84.2 85.7 91.5 72.5 82.5 92.0
GLoRIA[26] 84.1 86.8 89.1 85.1 88.5 92.1 66.5 80.5 88.0
BioViL[4] 82.0 85.4 88.6 79.8 81.6 90.5 - - -
LoVT [44] 85.1 86.5 89.3 85.5 88.5 92.2 - - -
MGCA [58] 85.8 87.7 89.2 86.1 89.6 92.0 74.8 84.8 92.3
PRIOR [11] 85.7 87.1 89.2 87.2 89.1 92.3 - - -
@@
ConVIRT[69] 46.5 53.9 53.8 20.0 45.4 35.5
GLoRIA[26] 46.7 56.4 55.0 51.8 59.5 58.9
BioViL[4] 47.3 57.7 55.6 54.6 64.3 62.8
MedCLIP [62] 47.6 58.0 55.9 56.3 69.9 66.7
MGCA [58] 47.1 57.4 55.4 53.1 61.9 61.1
REFERS [71] 52.4 59.9 58.6 60.6 71.9 69.0
CXR-CLIP [64] 51.8 61.2 58.5 60.2 69.2 64.6
@@ Methods
ConVIRT[69] 43.8 35.2 77.4
GLoRIA[26] 47.5 45.0 68.3
BioViL[4] 48.5 42.2 77.1
MedCLIP[62] 47.1 41.1 81.8
MGCA [58] 48.0 40.9 76.2
REFERS [71] 49.5 41.8 78.0
CXR-CLIP [64] 49.7 35.9 76.9
@@ 4 Experiments
We perform the pre-training on the current largest multi-modal medical dataset, MIMIC-CXR v2.0.0 [30] and evaluate our model on various downstream tasks, followed by an ablation study for probing purpose, which shows the proposed method’s superiority.
@@ 4.1 Implementation details
We employed a BERT model as the primary network for the image-text encoder and utilized ViT-g as the pre-trained ViT. The input image resolution was set to $224\times 224$, with a maximum text length of 95 tokens, and 32 learnable embeddings. The top and bottom codebook size were both set to 512 with feature dimension of 768. For optimization, we applied the AdamW optimizer [43] with parameters $\beta_{1}=0.9$, $\beta_{2}=0.95$, and a weight decay of 0.05. A cosine learning rate decay schedule was used, with a peak learning rate of 1e-4. We incorporated a warm-up phase for the initial 5% of training steps, starting with a learning rate of 1e-5. The pre-training process was conducted on four NVIDIA A100 GPUs.The implementation details are further elaborated in appendix 8.
@@ 4.2 Medical Vision-and-Language Benchmark
To assess the effectiveness of our method, we conducted experiments across three types of tasks: uni-modal, cross-modal, and multi-modal. We conducted experiments on MedUnifier with/without TIG and compared with previous studies as our main experimental results. All benchmark datasets used in experiments pertain specifically to radiology with detailed descriptions in the appendix 9. For baselines, we took reference from papers and implemented them using official codes: ConVIRT[69], GLoRIA[26], BioViL[4], LoVT [44], MedCLIP[62], MGCA [58], PRIOR [11], REFERS [71], CXR-CLIP [64], ViLT[2], R2Gen [9], PTUnifier [10], DCL [38], MOTOR [41], UniXGer[32], RoentGen[6], LLM-CXR[33].
Uni-modal tasks     assess the learned visual representations for image modality by applying uni-modal masking within classification scope on datasets such as RSNA Pneumonia [53], SIIM-ACR [66], and COVIDx [59]. To examine the model’s data efficiency, we fine-tune it using different proportions of the training data (1%, 10%, or 100%).
@@
require models to align vision and language modalities. We conduct experiments across three tasks: image-to-text retrieval (ITR), text-to-image retrieval (TIR), and zero-shot image classification (ZS). For ITR and TIR, we report cross-modal information retrieval metrics, the mean Average Precision at K (mAP@K), using the MIMIC 5x200 dataset. In the ZS task, we employ MIMIC 5x200, CheXpert 5x200 [26, 11], and draw 500 positive and 500 negative samples from the full RSNA Pneumonia dataset [53] for evaluation purpose.
@@
generate uni-modal content through multi-modal interaction. We carry out two kinds of experiments including image-grounded text generation (ITG) and text-grounded image generation (TIG). For ITG, we adopt the MIMIC-CXR held-out test set to evaluate the quality of generated reports. Standard natural language generation (NLG) criteria are used to assess the performance, including BLEUn [46], METEOR [12], and ROUGE-L [42]. For TIG, following [50], we first train two Pixelsnail models [8] to model multi-modal priors $\boldsymbol{z}_{q}^{top},\boldsymbol{z}_{q}^{bot}$. Then we sample latent encodings from both priors and generate new medical images by decoding latent encodings. For quantitative analysis, we present FID [23] scores.
@@ 4.3 Results and Analyses
To validate the effectiveness of MedUnifier, we conduct experiments on the above vision-and-language benchmark. The results of the main experiments are presented in Table 1,2, 3, 4 5 and Figure 6. We observe several noteworthy findings in our results. First, our model outperforms prior studies on uni-modal tasks across various downstream datasets, as shown in Table 1. This improvement suggests that integrating TIG significantly enhances the model’s ability to learn more transferable visual representations. Second, for cross-modal retrieval, MedUnifier model achieves the highest performance, demonstrating a superior ability to understand and integrate cross-modal content compared to other models (see Table 2). The obtained results further highlight the better performance of MedUnifier, suggesting that our approach supports the necessary complementary semantic data for cross-modal retrieval. We also observe that excluding the TIG module slightly decreases MedUnifier’s effectiveness in text-image retrieval tasks, which may be a result of an over-reliance on ITG that leads to an imbalance in vision-language fusion. In addition, MedUnifier demonstrates better performance on zero-shot classification tasks for both the MIMIC 5x200 and RSNA datasets in Table 3. However, GLoRIA outperforms our model on the CheXpert 5x200 dataset, likely due to its pre-training on the full CheXpert dataset with accompanying medical reports. Nonetheless, the results indicate the effectiveness of employing prompt ensembles within the proposed method, leading to enhanced overall performance improvements. Table 4 demonstrates that our models, both with and without the TIG module, surpass previous methods for image-grounded medical report generation. The Med-VLP framework also gains substantial advantages from incorporating causal language modelling, which together contributes to its improved performance. In Figure 3, we provide a comparison of the generated report and ground truth report. Finally, we conduct both quantitative and qualitative analyses on text-grounded image generation tasks in Table 5 and Figure 6 (see appendix). It highlights that MedUnifier with TIG achieve comparable performance for medical vision generation tasks with less model’s complexity (one-stage pre-training) then existing models [33, 32, 6], all of which require additional pre-training for image tokenizer or iterative denoising. From a direct visual inspection, the reconstructed visual samples are nearly indistinguishable from authentic radiographs. Moreover, synthetic samples generated from multi-modal priors demonstrate high diversity, highlighting their potential to augment out-of-distribution medical data effectively.
Figure 3: Comparison of ground truth and generated radiology reports reveals strong semantic alignment. In the top figure, both reports describe normal heart size, no pneumothorax or pleural effusion, and a normal cardiomediastinal silhouette, with the generated text adding details on osseous structures/intrathoracic processes. In the bottom figure, both reports align on pneumothorax and cardiomegaly. The same colours denote matched content between the generated sequences and the ground truth report.
@@
DCL[38] - - - 10.9 15.0 28.4
MOTOR[41] - 21.3 - 15.6 19.3 31.4
ViLT[2] - 21.3 - 9.2 - 29.6
R2Gen [9] 35.3 21.8 14.5 10.3 14.2 27.7
PTUnifier [10] - - - 10.7 - -
@@
Methods FID $\downarrow$ Methods FID $\downarrow$
UniXGer[32] 78.2 Validation set 17.2
RoentGen[6] 42.4 Reconstruction 27.2
LLM-CXR[33] 22.8 MedUnifier 46.2
Table 5: FID score [23] on reconstructed and synthetic images, using the MIMIC-CXR pre-training dataset as the reference dataset. The FID score using MIMIC-CXR validation is also provided for a fair comparison. Reconstructed images are generated by passing the training set through our MedUnifier. Synthetic images are produced from sampled and decoded latent encodings using the trained Pixelsnail models and VAE decoder.
ID Learning Objectives ITR Zero-shot cls Fine-tuned cls
ITC ITM ITG TIG mAP@1
@@
RSNA
SIIM
COVID
1 53.7 41.4 87.0 89.4 90.8
2 55.2 44.3 87.1 89.8 91.5
3 57.4 44.8 88.1 92.3 92.8
4 58.5 46.2 89.1 92.6 91.3
5 60.7 50.4 91.7 94.8 93.5
Table 6: Ablation studies on the different modules. The best performance is achieved using all objectives. Details see appendix 11.